OpenStudy (anonymous):

Sided limits at 5 of g(x). g(x) will be posted below.

OpenStudy (anonymous):

\[\frac{ x^2+x-30 }{ \left| x-5 \right| }\]

OpenStudy (amistre64):

might help to work the division, pretty sure the top has the bottom as a factor

OpenStudy (anonymous):

can you do that if it is absolute value on bottom?

OpenStudy (amistre64):

i cnt see why not since x<5 is negative, and x>5 is positive .... just have to keep that in mind as you develop a piecewise setup maybe?

OpenStudy (amistre64):

or let x = h+5 and work it for h to 0 .... just a random thought

OpenStudy (anonymous):

alright cool i got 11 and -11 and that was right thank you

OpenStudy (amistre64):

yay!! lol

OpenStudy (anonymous):

haha